Aparticle starts from rest at the point x=10 and moves along the x-axis with acceleration

function a(t) = 12t. Find the resulting position function x(t).


1
Expert's answer
2021-03-10T17:19:28-0500

Acceleration is the derivative of velocity with respect to time:


"a(t)=\dfrac{dv}{dt}."

We can find the velocity of the particle by integrating the acceleration function over time:


"\displaystyle\intop_{v_0}^v dv=\displaystyle\intop_{0}^t a(t)dt,""v-v_0=12\displaystyle\intop_{0}^t tdt,""v=v_0+6t^2=0+6t^2=6t^2."


Velocity is the deriative of position with respect to time:


"v(t)=\dfrac{dx}{dt}."

We can find the position of the particle by integrating the velocity function over time:


"\displaystyle\intop_{x_0}^x dx=\displaystyle\intop_{0}^t v(t)dt,""x-x_0=6\displaystyle\intop_{0}^t t^2dt,""x=x_0+2t^3,""x=10+2t^3"